2) Projectile through a tube: A projectile is fired horizontally with speed v, from the top of a cliff of hcight h. It immediately enters a horizontal tube of length x. There is friction between the projectile and the tube such that the projectile loses velocity with contant acceleration -a. After leaving the tube, the projectile undergoes normal projectile motion to the ground a. What is the total horizontal distance l traveled by the projectile measured from the base of the cliff expressed in terms of x, h, b. What value of x vields the maximum value of 1?
Answer #1

The projectile with initial horizontal velocity v_0 travels through the tube of length x with an acceleration -a to reach its final velocity at the end of the tube v=\sqrt{v_0^2-2ax} ( using v^2-u^2=2as )

With this horizontal velocity  v the projectile falls from hill of height h .

Horizontal distance travelled by projectile from the edge of tube to ground is (l-x) in time t .

horizontal distance = horizontal velocity x time taken to reach ground

(l-x)=vt ----(1)

Vertical distance travelled by projectile from the edge of tube to ground in time t is h

Using s=ut+\frac{1}{2}at^2 , vertical distance  h=0+\frac{1}{2}gt^2 ( initial vertical velocity = zero)

t=\sqrt{\frac{2h}{g}}

equation (1) can be written as (l-x)=\sqrt{v_0^2-2ax}\sqrt{\frac{2h}{g}}

That is \boxed{l=x+\sqrt{v_0^2-2ax}\sqrt{\frac{2h}{g}}}

b) For the maximum value of l , \frac{dl}{dx}=0

\frac{d}{dx}\left ( x+\sqrt{v_0^2-2ax}\sqrt{\frac{2h}{g}} \right )=0

1+\frac{1}{2}(v_0^2-2ax)^{-1/2}(-2a)\sqrt{\frac{2h}{g}}=0

(v_0^2-2ax)^{-1/2}=\frac{1}{a}\sqrt{\frac{g}{2h}}

(v_0^2-2ax)=a^2\,\,{\frac{2h}{g}}

\boxed{x=\frac{v_0^2}{2a}-\,\,{\frac{ah}{g}}}